OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 3x + 2y = 8 2x + 3y = 2 A.)4 B.)−2 C.)x can be any number as there are infinitely many solutions to this system D.)There is no x value as there is no solution to this system

OpenStudy (anonymous):

..?

OpenStudy (anonymous):

3x + 2y = 8 -----(1) 2x + 3y = 2 -----(2) Multiply equation (1) by 2 and equation (2) by 3 6x+4y=16 6x+9y=6 subtract -5y = 10 y=-2 plug this y into equation (1) 3x+2(-2) = 8 3x-4=8 3x=12 x=4 Im not sure if this is right, but thats my answer..

OpenStudy (shane_b):

You should end up with x=4 and y = -2

OpenStudy (anonymous):

Thank You! :)